Transverse waves with a speed of 50.0 m/s are to be produced on a stretched string having a

tension of 30.0 N. If this string has a total mass of 0.060 kg, (a) what is the total length of this

string? (b) How many times will the speed increases when the tension of the string is increased

three times?


1
Expert's answer
2021-02-03T16:12:05-0500

Answer

a) total mass is given as

"M=\frac{m}{L}=\frac{0.06}{5}=0.012" kg/m

b) relationship between velocity and force is given by

"v_1\propto\sqrt{F_1}"

When we increase tension at three time then

Velocity becomes

"v_2\propto\sqrt{3F_1}=1.732v_1"